Vector has magnitude of 20 km with quadrant bearing S 30° W and vector has magnitude of 35 km with quadrant bearing N 67°W. 




1
Expert's answer
2021-05-27T18:04:34-0400
"Solution"

The angle between y-axis and "12N" Vectors is ; Therefore

"\implies\ 360^\circ-\ 280^\circ\\
=80^\circ\\
Angle\ between\ vectors\\
\implies\theta=80+65\\
=\theta=145^\circ\\
Then\ using\ a\ parallelogram\\
Angle\ between\ vector\ \implies\ 180-\theta\\
\implies180-145\\
\implies\ 35^\circ\ 
Now,\ find\ the\ Resultant\ R\\
=\sqrt{(12)^2+(40)^2+2*12*40*Cos(35^\circ)}\\
=\sqrt{2530.38596}\\
=50.30\\
Direction(Use\ Sine)\\
=\frac{Sin(35^\circ)}{Sin\ \theta}=\frac{R}{40}\\
Sin\ \theta=\frac{40*Sin(35^\circ)}{50.30^\circ}\\
\theta=27.14^\circ"